# From LLMs to LLM-based Agents for Software Engineering: A Survey of Current, Challenges and Future

### v1

@@ V-B LLM-based Agents Tasks
LLM-based agents have shown significant potential and advantages by substantially improving task efficiency and effectiveness through multi-agent collaboration. Unlike traditional LLMs, LLM-based agents adopt a division of labor approach, breaking down complex tasks into multiple subtasks handled by specialized agents, this method can enhance task efficiency and improves the quality and accuracy of generated code to mitigate the hallucination from the single LLM.
@@
In [76], researchers proposed a self-collaboration framework where multiple ChatGPT (GPT-3.5-turbo) agents act as different roles to collaboratively handle complex code generation tasks. Specifically, the introduction of Software Development Methodology (SDM) divides the development process into three stages: analysis, coding, and testing. Each stage is managed by specific roles, and after completing their tasks, each role provides feedback and collaborates with others to improve the quality of the generated code. Experiment shows that this self-collaboration framework significantly improves performance on both the HumanEval and MBPP benchmarks, with the highest improvement reaching 29.9% in HummanEval compared to the SOTA model GPT-4. This result demonstrating the potential of collaborative teams in complex code generation tasks. Although it lacks external tool integration and dynamic adjustment capabilities, this framework exhibits common characteristics of LLM-based agents, such as role distribution, self-improvement ability, and excellent autonomous decision-making, these combined capabilities qualify it to be considered an LLM-based agent. Similarly, In [77], the LCG framework improved code generation quality also through multi-agent collaboration and chain-of-thought techniques, once again demonstrating the effectiveness of multi-agent collaboration in the software development process.
@@
The limitations of context windows was not discussed in previous studies, this has been thoroughly explored in a 2024 by University of Cambridge team. In [78], researchers introduced the L2MAC framework, which dynamically manages memory and execution context through a multi-agent system to generate large codebases, and achieved SOTA performance in generating large codebases for system design tasks. The framework is primarily divided into the following components: the processor, which is responsible for the actual generation of task outputs; the Instruction Registry, which stores program prompts to solve user tasks; and the File Storage, which contains both final and intermediate outputs. The Control Unit periodically checks the outputs to ensure that the generated content is both syntactically and functionally correct. The researchers conducted multiple experiments and compared with many novel methods like GPT-4, Reflexion, and AutoGPT, achieving a Pass@1 score of 90.2% on the HumanEval benchmark, showcasing its superior performance in generating large-scale codebases.
@@
Recently, many studies have begun to use LLM-based agents to simulate real software development processes, the paper  [79] introduced the MetaGPT framework, which enhanced problem-solving capabilities through standard operating procedures (SOPs) encoded in multi-agent collaboration. The entire process of the multi-collaboration framework simulates the waterfall life-cycle of software development, with each agent playing different roles and collaborating to achieve the goal of automating software development. LLM-based agents have also shown strong ability in automated software development, [80] proposed a multi-GPT agent framework that automates tasks such as project planning, requirement engineering, software design, and debugging, illustrating the potential for automated software development. Similarly  [81] introduced the model called CodePori, which is a novel model designed to automate code generation for extensive and complex software projects based on natural language prompts. In [82] the AgentCoder framework collaborates with programmer agents, test design agents, and test execution agents to generate and optimize code, outperforming existing methods, achieved SOTA performance on the HumanEval-ET benchmark with pass@1 of 77.4% compared to the previous state-of-the-art result of 69.5%, this result showcasing the advantages of multi-agent systems in code generation and testing.
@@
The purpose of integrating LLMs into agents from many framework is to enhance the self-feedback and reflection capabilities of the entire agent system. Because the current open-source LLMsgenerally have much lower capabilities in this aspect compared to proprietary models, the emergence of LLM-based agents can help bridge the gap between open-source models and the advanced capabilities of proprietary systems like GPT-4. [83] introduced the OpenCodeInterpreter framework, which improved the accuracy of code generation models by integrating code generation, execution, and human feedback. Based on CodeLlama and DeepSeekCoder, this framework performed close to the GPT-4 Code Interpreter on the HumanEval and MBPP benchmarks. The abbility of using external tools or APIs is another significant advantage of LLM-based agents, [84] proposed the Toolformer model, which significantly enhanced task performance by learning to call APIs through self-supervision. The framework Based on GPT-J (6.7B parameters) achieved significant performance improvements across multiple benchmark tasks, demonstrating the possibility of LLM-based agent brought by the external tool, the diverse choice of tools and architectures, allowing LLMs to continuously learn new things and improve themselves. Similarly, [85] enhanced LLMs’ interaction with external APIs through the ToolLLM framework, outperforming Text-Davinci-003 and Claude-2 on the ToolBench and APIBench benchmarks and excelling in multi-tool instruction processing.
@@ V-C Analysis
The main differences between LLM-based agents and traditional LLMs in software development applications mainly focus on the efficiency and autonomy, particularly in task division and collaboration. Traditional LLMs typically use a single model to handle specific tasks, such as generating code from text and code completion. However, this approach has limitations when dealing with complex tasks, especially regarding context window restrictions and the need for continuous feedback. LLM-based agents handle different subtasks through collaboration with clear division of labor, thereby enhancing task efficiency and quality. For example, in a code generation task, one agent generates the initial code, another designs test cases, and a third executes tests and provides feedback, thus achieving iterative optimization. Through task division, multi-agent systems, and tool integration, LLM-based agents can tackle more complex and broader tasks, improving the quality and efficiency of code generation. This approach overcomes the limitations of traditional LLMs also provides new directions and ideas for future software development research and applications, to frees programmers from the boring test suite generation.
@@
Figure 5: Illustration of Comparison Framework Between LLM-based Agent and LLM in Code Generation and Software Development
In software engineering task handling, there are subtle differences between LLMs and LLM-based agents in terms of task focus, approach, complexity and scalability, automation level, and task management. LLMs primarily focus on enhancing the code generation capabilities of a single LLM, including debugging, precision, evaluation. These methods typically improve specific aspects of code generation or evaluation through a single model, concentrating on performance enhancement within existing constraints, such as context windows and single-task execution. In contrast, LLM-based agents emphasize handling more complex and broader tasks through the collaboration of multiple specialized LLMs or frameworks, integrating tool usage, iterative testing, and multi-agent coordination to optimize the whole development process and easily surpass the state-of-art model in common benchmarks. The emeergence of multi-agent systems also brings more possibilities, this system can imitate the real software developer to perform the scrum development. Figure. 5 utilize studies [77] and [75] showcase the differences between LLM-based agents and LLMs on the same code generation task. The LLM-based agents system are able to perform multi-agent collaboration and simulate the real scrum development team in the industry. In contrast the LLMs on the right are normally use multi-LLMs to analysis mistakes from the test cases, and refine the initial generated code, but they lack autonomy and efficiency, as the test cases are manually generated by humans.
@@ V-E Evaluation Metrics
Various evaluation metrics are used to assess the performance of LLMs and LLM-based agents in code generation and software development. These metrics measures the models’ performance in specific tasks and how they improve the code generation and software development process. Table VI includes the distribution of evaluation metrics cited in this paper, encompassing both LLMs and LLM-based agents.
@@
TABLE VI: Evaluation Metrics in Code Generation and Software Development
@@
code repositories and notebooks Acceptance Rate, P
ercentage of Code Typed,
@@ VI Autonomous Learning and Decision Making
In this process, LLMs and LLM-based agents bring numerous possibilities, following the development of NLP technology, a lot of achievements have been made in the application of LLMs in this field. These models can handle complex language tasks and also demonstrate powerful reasoning and decision-making abilities, the research on voting inference using multiple LLMs calls has revealed new methods for optimizing performance, with the frequently used method called majority vote [89], this improves the accuracy of inference systems and ensures the selection of the optimal possibility. Additionally, the performance of LLMs in tasks such as automated debugging and self-correction has enhanced the system’s autonomous learning capabilities, achieving efficient error identification and correction. At the same time, the application of LLM-based agents in autonomous learning and decision-making is also a novel but popular topic, these agents can perform complex reasoning and decision-making tasks with the help from the LLM, and also improve their adaptability in dynamic environments through continuous learning and optimization. In this context, we have collected nineteen research papers on LLM-based agents in this field. This survey will provides a general review of these studies, analyzing the specific applications and technical implementations in autonomous learning and decision making.
@@ VI-A LLMs Tasks
The API call for the LLMs is one common applications, often requiring continuous calls to enable the model to make judgments and inferences, but does continuously increasing the number of calls always improve performance? In [90], researchers explored the impact of increasing LLM calls on the performance of composite reasoning systems. Paper analyze the voting inference design systems, the result showed that there is a non-linear relationship between the number of LLM calls and system performance; performance improves initially with more calls but declines after reaching a certain threshold. This research provides a theoretical basis for optimizing LLM calls, helping to allocate resources reasonably in practical applications to achieve optimal performance. However, the performance of Voting Inference Systems shows a non-monotonic trend due to the diversity of query difficulties, and the continuously increasing cost also needs to be considered.
@@
Autonomous learning is also applied in bug fixing, where researchers hope LLMs can continuously learn to fix bugs and eventually identify human oversights or common errors. In [91], the SELF-DEBUGGING method was proposed, enabling LLMs to debug code by analyzing execution results and natural language explanations. This method significantly improved the accuracy and sample efficiency of code generation tasks especially for complex problems. Experimental results on the Spider and TransCoder benchmarks showed that the SELF-DEBUGGING method increase the model’s accuracy by 2-12% which demonstrates the potential of LLMs in autonomous learning to debug and correct any erros. Another similar study introduced the AutoSD (Automated Scientific Debugging) technique [92], which simulates the scientific debugging process through LLMs, generating explainable patched code. Researchers evaluated AutoSD’s capabilities from six aspects: feasibility, debugger ablation, language model change, developer benefit, developer acceptance, and qualitative analysis. Result have shown that AutoSD can generate effective patches and also improve developers’ accuracy in evaluating patched code by providing explanations, its explainability function makes it easier for developers to understand and accept automatically generated patches. Although the above two studies primarily focus on automated debugging techniques, the frameworks designed in these studies automatically determine the optimal repair solution based on the debugging results after collecting sufficient information, and provide specific code implementations, which demonstrated the capability of autonomous decision-making and learning.
@@
Since the rise of LLMs applied to various fields, one research direction has been the rational analysis of their creativity and the exploration of their potential for continuous learning, this creativity also highly determined by the decision making capability of the models. [93] analyzed the outputs of LLMs from the perspective of creativity theory, exploring their ability to generate creative content, the study used metrics such as value, novelty, and surprise, finding that current LLMs have limitations in generating combinatorial, exploratory, and transformative creativity. Although LLMs can generate high-quality creative content, further research and improvement are needed to achieve true creative breakthroughs. Additionally, innovative responses generated by LLMs may come with the possibility of hallucination, a long-standing issue for large language models. Despite many techniques to mitigate its downsides, it still cannot be entirely prevented. There are many interesting experiments in decision making, such as having LLMs act as judges to determine whether a person has committed a crime [94]. A familiar attempt is to have a primary LLM interact with other LLMs. [95] explored the effectiveness of using LLMs as judges to evaluate other LLM-driven chat assistants. The study validated the consistency of LLM judgements with human preferences through the MT-Bench and Chatbot Arena benchmarks, with results showing that GPT-4’s judgments were highly consistent with human judgments across various tasks. This research demonstrates the potential of LLMs in simulating human evaluation, providing new ideas for automated evaluation and optimization.

### v2

@@ V-B LLM-based Agents Tasks
Multi-Agent Role Division for Software Development. LLM-based agents have shown significant potential and advantages by substantially improving task efficiency and effectiveness through multi-agent collaboration. Unlike traditional LLMs, LLM-based agents adopt a division of labor approach, breaking down complex tasks into multiple subtasks handled by specialized agents, this method can enhance task efficiency and improves the quality and accuracy of generated code to mitigate the hallucination from the single LLM.
@@
In [101], researchers proposed a self-collaboration framework where multiple ChatGPT (GPT-3.5-turbo) agents act as different roles to collaboratively handle complex code generation tasks. Specifically, the introduction of Software Development Methodology (SDM) divides the development process into three stages: analysis, coding, and testing. Each stage is managed by specific roles, and after completing their tasks, each role provides feedback and collaborates with others to improve the quality of the generated code. Experiment shows that this self-collaboration framework significantly improves performance on both the HumanEval and MBPP benchmarks, with the highest improvement reaching 29.9% in HummanEval compared to the SOTA model GPT-4. This result demonstrating the potential of collaborative teams in complex code generation tasks. Although it lacks external tool integration and dynamic adjustment capabilities, this framework exhibits common characteristics of LLM-based agents, such as role distribution, self-improvement ability, and excellent autonomous decision-making, these combined capabilities qualify it to be considered an LLM-based agent. Similarly, In [102], the LCG framework improved code generation quality also through multi-agent collaboration and chain-of-thought techniques, once again demonstrating the effectiveness of multi-agent collaboration in the software development process. Extending the multi-agent paradigm further, recent research has introduced the MapCoder framework [103], specifically designed for competition-level code generation tasks. MapCoder employs four distinct agents—Retrieval, Planning, Coding, and Debugging—each performing specialized tasks reflective of human programming practices. The Retrieval agent autonomously identifies relevant prior solutions, while the Planning agent generates step-by-step solution strategies. Subsequently, the Coding agent translates these strategies into executable code, and the Debugging agent iteratively refines this code based solely on sample input-output pairs, eliminating dependency on externally generated test cases. This robust division of responsibilities among specialized agents highlights the strength of multi-agent frameworks in significantly enhancing both code quality and problem-solving efficiency in complex programming environments.
@@
Overcoming Context Limitations for Large Codebases. The limitations of context windows were not discussed in previous studies, this has been thoroughly explored in 2024 by the University of Cambridge team. In [104], researchers introduced the L2MAC framework, which dynamically manages memory and execution context through a multi-agent system to generate large codebases, and achieved SOTA performance in generating large codebases for system design tasks. The framework is primarily divided into the following components: the processor, which is responsible for the actual generation of task outputs; the Instruction Registry, which stores program prompts to solve user tasks; and the File Storage, which contains both final and intermediate outputs. The Control Unit periodically checks the outputs to ensure that the generated content is both syntactically and functionally correct. The researchers conducted multiple experiments and compared with many novel methods like GPT-4, Reflexion, and AutoGPT, achieving a Pass@1 score of 90.2% on the HumanEval benchmark. Addressing similar scalability challenges, the Self-Organized multi-Agent framework (SoA) [105] proposes a novel solution based on dynamically scalable agent collaboration. Unlike traditional single-agent models constrained by limited context length, SoA employs multiple self-organized agents, each independently responsible for generating and refining smaller code components. This design allows the framework to automatically scale the number of agents in response to increasing complexity, thus maintaining manageable workloads per agent regardless of total codebase size. Evaluated against strong single-agent baselines like Reflexion [59], SoA achieves a significant improvement in Pass@1 accuracy, surpassing Reflexion by 5% on the HumanEval benchmark. The decentralized, scalable approach of SoA illustrates an effective strategy for overcoming single-agent context limitations, particularly beneficial in ultra-large-scale software development.
@@
Simulating the Software Engineering Lifecycle. Recently, many studies have begun to use LLM-based agents to simulate real software development processes, the paper  [57] introduced the MetaGPT framework, which enhanced problem-solving capabilities through standard operating procedures (SOPs) encoded in multi-agent collaboration. The entire process of the multi-collaboration framework simulates the waterfall life-cycle of software development, with each agent playing different roles and collaborating to achieve the goal of automating software development. LLM-based agents have also shown strong ability in automated software development, [27] proposed a multi-GPT agent framework that automates tasks such as project planning, requirement engineering, software design, and debugging, illustrating the potential for automated software development. Similarly  [106] introduced the model called CodePori, which is a novel model designed to automate code generation for extensive and complex software projects based on natural language prompts. In [107] the AgentCoder framework collaborates with programmer agents, test design agents, and test execution agents to generate and optimize code, outperforming existing methods, achieved SOTA performance on the HumanEval-ET benchmark with pass@1 of 77.4% compared to the previous state-of-the-art result of 69.5%, this result showcasing the advantages of multi-agent systems in code generation and testing. Building upon these foundations, more recent work explores how LLM-based agents can go beyond fixed workflows to emulate agile, dynamic, and context-aware development practices. One such framework is AGILECoder [108], a dynamic multi-agent architecture that explicitly incorporates Agile roles—PM, SM, Developer, Senior Developer, and Tester—and iteratively drives software development through structured sprints. It introduces a Dynamic Code Graph Generator (DCGG), which tracks cross-file dependencies to enhance context awareness for code generation, test case design, and debugging. Compared to waterfall-style pipelines like MetaGPT [57] or ChatDev [109], AGILECoder enables fine-grained planning, adaptive role-based review, and execution-based validation. Experiments on HumanEval, MBPP, and a new ProjectDev benchmark show that AGILECoder outperforms previous multi-agent systems in both pass@1 and executability, especially in large-scale, repository-level development tasks. This result demonstrates the importance of agile coordination and graph-based memory in driving scalable, real-world agent collaboration.Complementing this approach, a lightweight multi-agent framework proposed by [110] simplifies agile software development by assigning agents to handle user stories, task planning, coding, and pull request submissions in a structured yet less resource-intensive manner. Although it handles lower-complexity tasks without external tools or complex memory management, it reinforces the practicality of Agile-based multi-agent coordination for reducing manual interventions and maintaining coherent software outputs. These studies illustrate the versatility and scalability potential of LLM-based agents when explicitly aligned with Agile development methodologies.
@@
Enhancing Feedback, Tool Use, and Open-Source Model Performance. The purpose of integrating LLMs into agents from many framework is to enhance the self-feedback and reflection capabilities of the entire agent system. Because the current open-source LLMs generally have much lower capabilities in this aspect compared to proprietary models, the emergence of LLM-based agents can help bridge the gap between open-source models and the advanced capabilities of proprietary systems like GPT-4. [111] introduced the OpenCodeInterpreter framework, which improved the accuracy of code generation models by integrating code generation, execution, and human feedback. Based on CodeLlama and DeepSeekCoder, this framework performed close to the GPT-4 Code Interpreter on the HumanEval and MBPP benchmarks. The abbility of using external tools or APIs is another significant advantage of LLM-based agents, [112] proposed the Toolformer model, which significantly enhanced task performance by learning to call APIs through self-supervision. The framework Based on GPT-J (6.7B parameters) achieved significant performance improvements across multiple benchmark tasks, demonstrating the possibility of LLM-based agent brought by the external tool, the diverse choice of tools and architectures, allowing LLMs to continuously learn new things and improve themselves. Similarly, [113] enhanced LLMs’ interaction with external APIs through the ToolLLM framework, outperforming Text-Davinci-003 and Claude-2 on the ToolBench and APIBench benchmarks and excelling in multi-tool instruction processing.
Extending this paradigm, CodeAgent [54] addressed complex real-world, repository-level coding challenges through extensive integration of external programming tools, including retrieval systems, documentation parsing, and symbol navigation tools. By strategically orchestrating tool usage via multiple agent strategies (e.g., ReAct and Tool-Planning), CodeAgent significantly enhanced context-awareness and practical usability. Evaluations on the CODEAGENTBENCH benchmark demonstrated superior performance over commercial tools like GitHub Copilot, emphasizing the effectiveness of structured tool integration for tackling intricate software engineering tasks. A further advancement in this domain is provided by SWE-agent [53], which introduces a dedicated agent-computer interface (ACI) designed explicitly for LM-agents interacting with software engineering tasks. Unlike traditional Linux shells or general-purpose interfaces, the ACI provides simplified, LM-friendly commands, immediate feedback, and built-in guardrails for common software tasks such as file navigation, code editing, and test execution. This carefully tailored interface significantly enhances LLM-agent efficiency, achieving state-of-the-art pass@1 rates on SWE-bench (12.5%) and HumanEvalFix (87.7%), clearly demonstrating that interface design is critical in enabling effective, autonomous software engineering by LLM-based agents. The CodeAct framework proposed in [114] allows agents to generate Python code directly executable by a Python interpreter, enabling agents to dynamically revise actions and integrate existing Python libraries seamlessly. CodeAct’s structured executable actions allow LLMs to effectively utilize control flow mechanisms such as loops and conditionals to handle complex multi-step tool invocations within a unified action framework. This framework underline the potential of executable code actions in bridging the gap between open-source and proprietary models by capitalizing on LLMs’ inherent code comprehension skills, thus enhancing their practical usability in diverse, complex software engineering tasks.
Interactive Requirement Clarification and Self-Planning. ClarifyGPT [115] is a novel LLM-based agent framework designed to address the ambiguity in user-provided natural language requirements during code generation. Unlike conventional prompting methods that directly generate code regardless of input clarity, ClarifyGPT first determines whether a requirement is ambiguous via a two-step code consistency check, which samples multiple code completions and evaluates their behavioral divergence on synthesized test inputs. If inconsistency is detected, the system engages in a reasoning-driven prompt strategy to generate targeted clarifying questions. Upon receiving responses (either from users or a high-fidelity simulator), the agent refines the original requirement and re-prompts the LLM to generate code. This pipeline, involving perception, planning, action, and reflection, exemplifies typical LLM-based agent characteristics. Evaluation on five benchmarks (MBPP, HumanEval, CoderEval, etc.) demonstrates that ClarifyGPT significantly improves the Pass@1 accuracy, achieving a relative gain of 11.66% with GPT-4 and 15.00% with ChatGPT. These improvements are particularly notable on ambiguous tasks, confirming the effectiveness of proactive clarification in agent-driven development workflows.
@@ V-C Analysis
The evolution from traditional LLM-based approaches to more autonomous LLM-based agent frameworks in software development underscores a fundamental shift in how code generation is conceptualized, orchestrated, and evaluated. Traditional LLM methods typically center on improving a single model’s ability to generate, debug, or refine code. While powerful, these methods often face context window constraints, limited memory, and a linear development workflow that depends heavily on human-provided test cases or clarifications. In contrast, LLM-based agent systems distribute responsibilities across multiple specialized agents (or roles) and often integrate external tools to emulate more realistic and adaptive software engineering practices.
One key difference is in the degree of autonomy and collaboration. Traditional LLMs, even those enhanced with techniques like few-shot prompting [37], chain-of-thought [16], or feedback-driven refinement [116], tend to operate as a single model with minimal division of labor. They excel in generating snippets or providing incremental improvements, but often struggle in orchestrating an entire development process, especially when large codebases or complex cross-file dependencies are involved. By contrast, the newer LLM-based agent frameworks discussed in Section V-B (e.g., SoA, L2MAC, AGILECoder, CodeAgent) simulate multi-step development pipelines through distributed roles or agents for planning, coding, testing, and debugging. This division of labor allows for context expansion beyond a single model’s window limit, as well as iterative feedback and refinement that better aligns with real-world team-based software development processes.
Another prominent distinction lies in tool integration. Although LLMs can incorporate limited forms of external knowledges, such as print debugging [92], test case execution [97, 98], or retrieval mechanisms [99], they rarely exhibit flexibility. In agent systems, each specialized role (e.g., a “retrieval agent” or a “debug agent”) can invoke dedicated libraries, file navigation interfaces, or testing harnesses without human intervention. This substantially boosts effectiveness in tasks like repository-level code completion, where scanning large amounts of context or retrieving relevant snippets from extensive codebases becomes essential. Moreover, it helps address the gap between open-source LLMs and advanced proprietary systems by enhancing self-reflection and externally verifiable feedback mechanisms.
Scalability and life-cycle simulation further differentiate the two paradigms. Traditional LLM-based methods do handle multi-turn queries, such as iterative code completion or user-guided debugging, but they do not generally organize these interactions under a coherent, larger software development methodology. In contrast, many agentic systems like MetaGPT [117], ChatDev [109] and AGILECoder [108] explicitly model entire development cycles, whether it is the classic waterfall model or agile “sprints”. By dynamically coordinating specialized agents, these frameworks emulate real-world team strategies: clarifying requirements, planning sprints, writing code, and designing or running tests. As a result, tasks that go beyond single-file code generation (like orchestrating multiple modules, dealing with cross-dependencies, or generating large-scale projects) become more tractable.
@@
Figure 6: Illustration of Comparison Framework Between LLM-based Agent and LLM in Code Generation and Software Development.
As illustrated in Figure 6 (adapted from [102] and [96]), the left-hand flow represents an LLM-based agent system that divides software tasks among specialized roles simulating an industry Scrum team. These roles collaboratively implement, test, and refine code through iterative coordination. In contrast, the right-hand flow depicts how traditional LLMs might process multiple prompts or code analyses in isolation, lacking the continuous autonomous coordination and tool-driven synergy inherent to multi-agent environments. While single-model approaches can achieve high pass@k performance on established benchmarks, the agent-based paradigm demonstrates broader coverage of large-scale, complex tasks, frequently surpassing state-of-the-art results.
This shift toward LLM-based agents reflects the growing demand for holistic, end-to-end solutions in software development. Traditional LLMs excel at isolated tasks but face limitations due to constrained context awareness, insufficient feedback mechanisms, and minimal autonomy. In contrast, LLM-based agents leverage role specialization, iterative refinement, tool integration, and dynamic context management. These capabilities address critical performance gaps in large-scale, multi-file coding challenges and enable autonomous workflows that increasingly mirror human-led engineering processes.
@@ V-E Evaluation Metrics
Various evaluation metrics are used to assess the performance of LLMs and LLM-based agents in code generation and software development. These metrics measures the models’ performance in specific tasks and how they improve the code generation and software development process. Table VII includes the distribution of evaluation metrics cited in this paper, encompassing both LLMs and LLM-based agents.
@@
TABLE VII: Evaluation Metrics in Code Generation and Software Development
@@
 [97] MBPP, HumanEval pass@k, pass@1@m No
 [98] CodeContests pass@k No
 [99] None (real-world prompts + OSCAT lib) Manual simulation and correctness
via OpenPLC No
 [100] CodeContests, LeetCode (recent) n@k (1@k, 2@k, 10@k, 100@k) No
@@
 [115] MBPP-sanitized, MBPP-ET, HumanEval/-ET, CoderEval Pass@1 Yes
 [108] HumanEval, MBPP, ProjectDev Pass@1, Executability Rate, #Errors, Runtime Stats Yes
 [110] Real-world case studies Human Review, Task Success Rate Yes
 [105] HumanEval Pass@1 Yes
 [53] SWE-bench, SWE-bench Lite, HumanEvalFix Pass@1, %Resolved, pass@k, \$Avg. Cost Yes
 [114] API-Bank, M3ToolEval Success Rate, Avg. Number of Turns, Tool Correctness Yes
 [103] HumanEval/-ET, EvalPlus, MBPP/-ET,
APPS, xCodeEval, CodeContests Pass@k Yes
 [54] CODEAGENTBENCH, HumanEval Pass@1 Yes
@@
code repositories and notebooks Acceptance Rate, Percentage
of Code Typed,
@@ VI Autonomous Learning and Decision Making
In this process, LLMs and LLM-based agents bring numerous possibilities, following the development of NLP technology, a lot of achievements have been made in the application of LLMs in this field. These models can handle complex language tasks and also demonstrate powerful reasoning and decision-making abilities, the research on voting inference using multiple LLMs calls has revealed new methods for optimizing performance, with the frequently used method called majority vote [119], this improves the accuracy of inference systems and ensures the selection of the optimal possibility. Additionally, the performance of LLMs in tasks such as automated debugging and self-correction has enhanced the system’s autonomous learning capabilities, achieving efficient error identification and correction. At the same time, the application of LLM-based agents in autonomous learning and decision-making is also a novel but popular topic, these agents can perform complex reasoning and decision-making tasks with the help from the LLM, and also improve their adaptability in dynamic environments through continuous learning and optimization. In this context, we have collected 23 research papers on LLM-based agents in this field. This survey will provides a general review of these studies, analyzing the specific applications and technical implementations in autonomous learning and decision making.
@@ VI-A LLMs Tasks
API Utilization and Optimization in LLM Workflows. The API call for the LLMs is one common applications, often requiring continuous calls to enable the model to make judgments and inferences, but does continuously increasing the number of calls always improve performance? In [120], researchers explored the impact of increasing LLM calls on the performance of composite reasoning systems. Paper analyze the voting inference design systems, the result showed that there is a non-linear relationship between the number of LLM calls and system performance; performance improves initially with more calls but declines after reaching a certain threshold. This research provides a theoretical basis for optimizing LLM calls, helping to allocate resources reasonably in practical applications to achieve optimal performance. However, the performance of Voting Inference Systems shows a non-monotonic trend due to the diversity of query difficulties, and the continuously increasing cost also needs to be considered.
@@
Autonomous learning Learning from Feedback. Autonomous learning is also applied in bug fixing, where researchers hope LLMs can continuously learn to fix bugs and eventually identify human oversights or common errors. In [121], the SELF-DEBUGGING method was proposed, enabling LLMs to debug code by analyzing execution results and natural language explanations. This method significantly improved the accuracy and sample efficiency of code generation tasks especially for complex problems. Experimental results on the Spider and TransCoder benchmarks showed that the SELF-DEBUGGING method increase the model’s accuracy by 2-12% which demonstrates the potential of LLMs in autonomous learning to debug and correct any erros. Another similar study introduced the AutoSD (Automated Scientific Debugging) technique [122], which simulates the scientific debugging process through LLMs, generating explainable patched code. Researchers evaluated AutoSD’s capabilities from six aspects: feasibility, debugger ablation, language model change, developer benefit, developer acceptance, and qualitative analysis. Result have shown that AutoSD can generate effective patches and also improve developers’ accuracy in evaluating patched code by providing explanations, its explainability function makes it easier for developers to understand and accept automatically generated patches. Although the above two studies primarily focus on automated debugging techniques, the frameworks designed in these studies automatically determine the optimal repair solution based on the debugging results after collecting sufficient information, and provide specific code implementations, which demonstrated the capability of autonomous decision-making and learning. Another framework ALGO [100], which was previously discussed under the theme of code generation and software development due to its significant improvements in program synthesis. However, ALGO is also highly relevant to autonomous learning and decision-making. It embodies a verifier-guided learning paradigm in which the LLM autonomously generates brute-force oracles, constructs test cases, and iteratively compares outputs to verify correctness and guide efficient code synthesis. This self-supervised refinement loop allows the model to make informed decisions about code quality and correctness without external supervision, exemplifying autonomous reasoning and learning in practice.
@@
LLM-Based Decision Making and Evaluation. Since the rise of LLMs applied to various fields, one research direction has been the rational analysis of their creativity and the exploration of their potential for continuous learning, this creativity also highly determined by the decision making capability of the models. [123] analyzed the outputs of LLMs from the perspective of creativity theory, exploring their ability to generate creative content, the study used metrics such as value, novelty, and surprise, finding that current LLMs have limitations in generating combinatorial, exploratory, and transformative creativity. Although LLMs can generate high-quality creative content, further research and improvement are needed to achieve true creative breakthroughs. Additionally, innovative responses generated by LLMs may come with the possibility of hallucination, a long-standing issue for large language models. Despite many techniques to mitigate its downsides, it still cannot be entirely prevented. There are many interesting experiments in decision making, such as having LLMs act as judges to determine whether a person has committed a crime [124]. A familiar attempt is to have a primary LLM interact with other LLMs. [125] explored the effectiveness of using LLMs as judges to evaluate other LLM-driven chat assistants. The study validated the consistency of LLM judgements with human preferences through the MT-Bench and Chatbot Arena benchmarks, with results showing that GPT-4’s judgments were highly consistent with human judgments across various tasks. This research demonstrates the potential of LLMs in simulating human evaluation, providing new ideas for automated evaluation and optimization.